OpenStudy (sparklyme):

Having trouble FACTORING this... n^2+21n = 0

OpenStudy (sparklyme):

@mathmate Are you available for some help?

jimthompson5910 (jim_thompson5910):

think of n^2 as n*n `n^2+21n = 0` is the same as `n*n+21*n = 0` we can divide both terms (n*n and 21*n) by the common factor n so we can pull it out In other words, n*n+21*n = n*(n+21) hopefully you can see that if you distribute the n back in, you get n*n + 21*n back again?

OpenStudy (sparklyme):

@jim_thompson5910 I'm still kind of confused...

jimthompson5910 (jim_thompson5910):

Let's say we replace n with a number. What number do you want to replace n with?

jimthompson5910 (jim_thompson5910):

You can pick any number

OpenStudy (sparklyme):

@jim_thompson5910 2

jimthompson5910 (jim_thompson5910):

so we would go from `n*n + 21*n` to `2*2 + 21*2` agreed?

OpenStudy (sparklyme):

yes

jimthompson5910 (jim_thompson5910):

the common factor here is 2 we can pull this out 2*2 + 21*2 = 2*(2+21) think of it as distribution but in reverse

jimthompson5910 (jim_thompson5910):

if n = 3, then we'd have 3*3 + 21*3 = 3*(3+21)

OpenStudy (sparklyme):

so in 2(2+21), 2 before the () would actually be n, correct?

jimthompson5910 (jim_thompson5910):

yes because n represents any number (not just 2)

OpenStudy (sparklyme):

ok. but how am i going to find n?

jimthompson5910 (jim_thompson5910):

same with the 2 inside the parentheiss

OpenStudy (sparklyme):

or am i not supposed to? because it just says to factor it.

jimthompson5910 (jim_thompson5910):

you leave n as itself. There's no need to actually find the value of n

OpenStudy (sparklyme):

oh. ok.

jimthompson5910 (jim_thompson5910):

you just go from `n^2+21n` to `n(n+21)`

jimthompson5910 (jim_thompson5910):

so n^2+21n = 0 is the same as n(n+21) = 0

OpenStudy (sparklyme):

right....so then it would be (-21+21)=0 so -21 would be the answer, correct?

jimthompson5910 (jim_thompson5910):

-21 is one solution since it is the solution to n+21 = 0

jimthompson5910 (jim_thompson5910):

n(n+21) = 0 means that n = 0 or n+21 = 0

jimthompson5910 (jim_thompson5910):

so if n(n+21) = 0, then n = 0 or n = -21

OpenStudy (sparklyme):

oh......so if there is x^2 (or another variable) + a number (lets say 32)x, then x would automatically equal 0?

jimthompson5910 (jim_thompson5910):

no, in your original equation, you have n^2+21 set equal to zero

jimthompson5910 (jim_thompson5910):

n^2+21n I mean

jimthompson5910 (jim_thompson5910):

remember that if A*B = 0 then either A = 0 or B = 0

OpenStudy (sparklyme):

well if it was set equal to zero so if it was like: x^2 + 32x +0 it would automatically equal 0?

OpenStudy (sparklyme):

i mean =0

jimthompson5910 (jim_thompson5910):

if you had x^2 + 32x = 0, then it would factor to x*(x+32) = 0

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):